OpenStudy (anonymous):

(2/3x)^3 help me simplify

OpenStudy (theviper):

\[\Large{\left( \frac{a}{b} \right)^3=\frac{a^3}{b^3}}\]

OpenStudy (theviper):

So can u do it now :)

OpenStudy (anonymous):

not sure

OpenStudy (theviper):

Ok do here i will see :)

OpenStudy (anonymous):

see here http://www.wolframalpha.com/input/?i=%282%2F3x%29%5E3

OpenStudy (theviper):

Hey @best.shakir this site do not allow anyone to give direct answers read http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

would it be a^6b^6

OpenStudy (theviper):

\[\Large{\left( \frac{2}{3x} \right)^3=\frac{(2)^3}{(3x)^3}=??}\]

OpenStudy (anonymous):

6/9

OpenStudy (theviper):

no 2 * 2 * 2 = 8 similarly 3x * 3x * 3x = 27x^3

OpenStudy (anonymous):

so it is 27x^3

OpenStudy (theviper):

So answer is :-\[\Huge{\color{green}{\frac{(2)^3}{(3x)^3}=\frac{8}{27x^3}}}\]

OpenStudy (theviper):

Now understood ??

OpenStudy (anonymous):

ok thanks

OpenStudy (theviper):

ur welcome :)